Test the claim that the mean GPA of night students is smaller than 2.8 at the .10 significance level. The null and alternative hypothesis would be: H1 : p < 0.7 H1ιμ>2.8 H1 : μ < 2.8 Ho:p 0.7 Ho:p 0.7 Ho: 2.8 The test is: left-tailed right-tailed two-tailed Based on a sample of 65 people, the sample mean GPA was 2.76 with a standard deviation of 0.05 The test statistic is: decimals) The critical value is: decimals) Based on this we to (to 2 OFail to reject the null hypothesis Reject the null hypothesis
Answer #1

Given that, sample size ( n ) = 65

sample mean = 2.76 and standard deviaton = 0.05

significance level = 0.10

The null and the alternative hypotheses are,

H_0 : \mu = 2.8

H_1: \mu < 2.8

Test statistic is,

Z = \frac { 2.76 - 2.8}{\frac { 0.05}{\sqrt {65}}}= -6.45

The test statistic is : -6.45

the critical value at 0.10 significance level is: -1.28

Here, test statistic = -6.45 < -1.28

so, we Reject the null hypothesis